The new owner of a student apartment building near campus was considering establishing a No Pets rule. He decided to survey the current residents to see how many owned pets and what kind of pets they owned. The building had 58 units. The previous owner had not allowed cats or dogs or other mammals, but the residents reported a variety of other pets: 30 owned birds, 38 owned fish, and 30 owned frogs, lizards, or snakes. Of those how had fish, 19 also had a frog, lizard, or snake.The owner found that 18 had both fish and birds and 15 kept a bird and a frog, lizard, or snake. Lastly, he determined that 10 owned fish, abird, lizard, or snake. How many residents could have reported no pets? How many residents owned a bird but not a fish? How many residents do not own a bird?

Question 1: How many residents owned a bird but no fish?

Question 2: How many residents did not own a bird?

I will rate**
Answer #1

Tetal 58 B A n - 3o atet dtg birds A owned et b+gt8 -38 B owned Lish 2 d+gtc 3 3 0 C = albo had og /izad snale. theu who haddig = 15 . dtlo 15 = 19 = lo + de 5 etg=18 e+lo= 18 e 8 2 frenm Fam 3 digtBtc = 30 + g+B = 38 e+ bt 5 + lo+9 +C30 b+ l0t 9=38n 2 Residents w No peto but net a bird Residonts ownid atd 12 kird. did nat own 3) Residants I+ 9t6+ 2 2.8